# Out-of-Distribution Semantic Occupancy Prediction

### v1

@@ 6 Conclusion
In this study, we look into out-of-distribution semantic occupancy prediction to address OoD voxel detection in 3D space. To address the scarcity of anomaly annotations in existing datasets, we developed a Synthetic Anomaly Integration Pipeline and introduced two datasets, VAA-KITTI and VAA-KITTI-360, for evaluating OoD detection in 3D occupancy tasks. We put forward OccOoD to integrate OoD detection into 3D semantic occupancy prediction. By combining voxel-based and BEV representations, OccOoD effectively captures both local details and global context, thereby enhancing its semantic and geometric understanding. Experiments show that OccOoD delivers state-of-the-art OoD detection performance on VAA-KITTI and VAA-KITTI-360, while also maintaining competitive occupancy prediction performance on public benchmarks.
Limitation: The performance of current semantic occupancy prediction methods is limited by geometric inaccuracies in estimated occupancy grids, which adversely affect precision of anomaly detection. We anticipate that future advancements in the field will substantially address these limitations, leading to more robust 3D perception.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. 62473139) and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B20).

### v2

@@ VI Conclusion
In this paper, we look into out-of-distribution semantic occupancy prediction to address OoD voxel detection in 3D space. To address the lack of anomaly annotations in existing datasets, we propose a Realistic Anomaly Augmentation and develop three evaluation datasets: VAA-KITTI, VAA-KITTI-360, and VAA-STU. Our method, OccOoD, integrates OoD detection into 3D semantic occupancy prediction by combining voxel-based and BEV representations, effectively capturing both local detail and global context for enhanced semantic and geometric reasoning. Experiments show that OccOoD achieves state-of-the-art OoD detection performance on these datasets while maintaining competitive occupancy prediction accuracy on standard benchmarks.